OpenStudy (anonymous):

Emily is converting 9 feet per minute into inches per minute. Which conversion factor should she use?

OpenStudy (accessdenied):

We start with feet per minute, and we are converting to inches per minute. So the minutes unit stays the same. We just need to convert inches to feet somehow. If you know how many inches are in a foot, we can get somewhere!

OpenStudy (accessdenied):

In order to make a conversion, we want to cancel out the unit we are converting by applying an inverse operation. \( \dfrac{\cancel{\text{feet}}}{\text{minute}} \times \dfrac{\text{inches}}{\cancel{\text{feet}}} \to \dfrac{\text{inches}}{\text{minute}} \)

OpenStudy (accessdenied):

Does this make sense thus far? :)